OpenStudy (anonymous):

find the vertex, the line of symmetry, the maximum or minimum value of the quadratic function.. f(x)=-2X^2+2X+7

OpenStudy (zarkon):

x=1/2

OpenStudy (anonymous):

i have to put it in as a fraction. like 1/2

OpenStudy (zarkon):

(1/2,15/2)

OpenStudy (anonymous):

yea i got 1/2 now then is it 15/2?

OpenStudy (anonymous):

wud the equation be x=15/2?

OpenStudy (zarkon):

x=1/2 is the line of symmetry and the vertex is at (1/2,15/2)

OpenStudy (anonymous):

yea i got those answers would my equation be correct?

OpenStudy (zarkon):

why would you need x=15/2?

OpenStudy (zarkon):

15/2 is the y-value that corresponds to the maximum.

OpenStudy (anonymous):

it is asking for the equation of the line of symmetry

OpenStudy (anonymous):

it would be x=1/2